# Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Spatial Reasoning Questions

### v3

@@ Abstract
Spatial reasoning remains challenging for Large Language Models (LLMs), which struggle with spatial data retrieval and reasoning. We propose Spatial Retrieval-Augmented Generation (Spatial-RAG), a framework that extends RAG to spatial tasks by integrating sparse spatial retrieval (spatial databases) and dense semantic retrieval (LLM-based similarity). A multi-objective ranking strategy balances spatial constraints and semantic relevance, while an LLM-guided generator ensures coherent responses. Experiments on a real-world tourism dataset show that Spatial-RAG significantly improves spatial question answering, bridging the gap between LLMs and spatial intelligence.
RAG, spatial, LLM agent, tourism question answering
@@ 1 Introduction
Spatial questioning has long been a fundamental domain (Obe & Hsu, 2021), which adeptly handles a variety of spatial questions, from identifying closest neighbors to detecting line-polygon intersections. However, traditional spatial questioning systems rely on specialized spatial query languages, which are vastly different from human language, making them inaccessible to layman users. More critically, these systems lack the ability to infer complex spatial and semantic relationships from the nuanced, context-rich nature of human text, limiting their applicability to real-world question-answering scenarios. Recent advancements in Large Language Models (LLMs) have transformed many fields within machine learning (ML), particularly in understanding and generating human-like text. This progress has inspired early efforts to bridge the gap between spatial questioning and natural language by directly eliciting spatial knowledge from LLMs. These efforts span a range of applications, including geographic encyclopedic question answering (Mai et al., 2021, 2020; Chen, 2014; Scheider et al., 2021; Contractor et al., 2021a), geolocalization (Zhou et al., 2024; Haas et al., 2024), and automated High-definition map generation (Li et al., 2022; Liu et al., 2023). Despite these advancements, recent studies indicate that LLMs exhibit significant limitations in spatial reasoning (Mai et al., 2024; Roberts et al., 2023), and struggle with even basic spatial tasks, such as geoparsing (Mai et al., 2024) and understanding relative spatial relationships (Majic et al., 2024). This gap becomes particularly evident when handling real-world spatial reasoning tasks, such as the one illustrated in Figure 1, where the system must recommend a restaurant along a specific route within a predefined area.
@@
Figure 1: An example of a real-world spatial reasoning question with nearby spatial objects. Areas that satisfy the spatial constraint are highlighted in blue.
@@
This paper tackles a novel challenge: augmenting LLMs with spatial reasoning capabilities. Specifically, we extend Retrieval-Augmented Generation (RAG) into spatial information retrieval and reasoning, bridging the gap between structured spatial databases and unstructured textual reasoning. RAG has demonstrated its effectiveness in knowledge-intensive tasks, such as question answering (QA) (Siriwardhana et al., 2023), by retrieving domain-specific documents to enhance LLM responses. However, existing RAG systems primarily focus on retrieving and generating textual content and lack the spatial intelligence required for spatial reasoning tasks, especially tasks that involve understanding and computing complex spatial relationships among geometries, including points, polylines, and polygons.
@@
Spatial questioning is intrinsically challenging, requiring a synergistic combination of text-based inference and spatial computations. For example, as illustrated in Figure 1, answering the question requires LLM to elicit and formulate the user’s textual request into the problem of ”finding points near the polyline” and solve it based on a spatial map (database). Then, it also requires inferring user preference to select the spatially and semantically preferred candidates. Thus, the system must seamlessly integrate structured spatial retrieval with unstructured text-based reasoning, ensuring both spatial accuracy and contextual understanding. To achieve this, we introduce Spatial Retrieval-Augmented Generation (Spatial-RAG), a novel framework that unifies text-guided spatial retrieval with spatially aware text generation. Specifically, to identify spatially relevant candidate answers, we propose a novel spatial hybrid retrieval module combining sparse and dense retrievers. To rank the candidates and generate the final answers, we propose to fuel the generator with retrieved results on the Pareto frontier based on a spatial and semantic joint ranking strategy. Our contributions are summarized as follows:
@@
  • A generic Spatial RAG framework: We introduce spatial-RAG, the first framework that extends RAG to spatial question answering, to tackle a broad spectrum of spatial reasoning tasks, such as geographic recommendation, spatially constrained search, and contextual route planning. Our approach seamlessly integrates spatial databases, LLMs, and retrieval-based augmentation, enabling effective handling of complex spatial reasoning questions directly within the familiar operational paradigm of LLMs.
@@
  • Sparse-dense spatial hybrid retriever: We propose a hybrid retrieval mechanism that combines sparse retrieval (SQL-based structured queries) with dense retrieval (LLM-powered semantic matching). This dual approach ensures that retrieved results align both spatially and semantically with the user’s query, significantly improving retrieval accuracy in spatial contexts.
@@
  • Multi-objective guided spatial text generator: To handle both spatial constraints and textual reasoning in the spatial question-answering task, we introduce a multi-objective optimization framework that dynamically balances trade-offs between spatial and semantic relevance. This ensures that generated responses are both geometrically accurate and linguistically coherent.
@@
  • Real-World Evaluation: We evaluated our method on a real-world dataset collected from a Tourism website that has user questions and reviews about different spatial entities. Experiments on this dataset reveal the ability to handle real-world spatial reasoning questions.
@@
Through these innovations, Spatial-RAG significantly enhances the spatial reasoning capabilities of LLMs, bridging the gap between structured spatial databases and natural language question answering.
@@ 2.1 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) is a hybrid approach that integrates retrieval systems and generative models to enhance factual accuracy and contextual relevance in natural language generation (Fan et al., 2024). Unlike conventional language models that rely solely on parametric memory, RAG dynamically retrieves relevant external knowledge before generating a response. One of the foundational works in RAG is Lewis et al. (2020), where a retrieval module fetches relevant passages from a large-scale knowledge corpus (e.g., Wikipedia), which are then fused with the question context to generate a more informed response. This technique has proven particularly effective in open-domain question answering (QA), fact verification, and context-aware text generation. RAG systems have expanded beyond text and document retrieval to incorporate a wide variety of data types (He et al., 2024) — tables, graphs, charts, and diagrams. While RAG has been widely explored, its application in spatial reasoning question answering remains an unexplored research area. Existing studies have primarily focused on knowledge-grounded dialogues (Yu et al., 2024) but often struggle with integrating spatial computation into the question-answering process effectively.
2.2 Spatial Questions
@@
Spatial questions in domain-specific applications can generally be categorized into two distinct types: 1) Textual Knowledge-based Spatial Questions These are spatial questions that can be answered by traditional QA methods without the need for spatial computation and reasoning (Liétard et al., 2021). For example, the question ”What is the population of Los Angeles city?” falls under this category. Despite their spatial context, these questions are essentially text-based and, hence, can be effectively addressed using traditional Retrieval-Augmented Generation (RAG) methods (Christmann & Weikum, 2024). 2) Spatial Reasoning Questions This category encapsulates spatial questions that demand a model’s capability to comprehend and reason with spatial data and spatial relationships. A common example is a model being presented with textual information describing the spatial relationships among multiple objects (Li et al., 2024). An example question could be, ”What is the position of object A relative to object B?”, where objects A and B are locations or entities specified on the map. Resolving such queries requires a profound understanding of spatial concepts and robust reasoning skills, which largely depend on the model’s training to handle spatial data. Several studies (Mai et al., 2024; Roberts et al., 2023) have investigated the capacity of LLMs to understand spatial concepts, yet these models often struggle with accurate reasoning even after fine-tuning. Other research (Li et al., 2023) has attempted to enhance this ability by converting geolocation coordinates into addresses to enrich the semantic context. However, these improvements tend to be marginal and are mostly limited to straightforward reasoning tasks like describing positions. Moreover, many existing methods rely on predefined sets of actions tailored to specific tasks.
@@ 3 Problem Formulation
In this study, our primary focus is Spatial Reasoning Questions. We formulate the problem as follows: Given a question $q$, the system aims to generate an answer $y$,
@@
s.t. $y\in C_{s}(q),\quad y\in C_{k}(q),$
$\lambda_{s}\geq 0,\quad\lambda_{k}\geq 0,$
$\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1,$
@@
To solve this problem, there are three questions to answer: 1) How to acquire the spatial candidate set $C_{s}$? 2) How to synergize the spatial and texts by evaluating $f_{s}(q,y)$? 3) How to trade off the spatial and semantic aspects?
4 Methodology
@@ 4.1 Overview
Our proposed framework, Spatial Retrieval-Augmented Generation (Spatial-RAG), is illustrated in Figure 2. Spatial-RAG consists of three key stages: First, to construct the spatial candidate set $C_{s}$, the system must precisely define spatial constraints and then retrieve spatial objects that satisfy them. As depicted in Figure 2 (Sparse Spatial Retrieval), we achieve this by parsing the input natural language questions into a spatial SQL query, which will be executed on the spatial database to efficiently retrieve relevant spatial objects from the database. This process is detailed in Section 4.2. Second, to effectively compute spatial relevance $f_{s}(q,y)$ while integrating textual information, we propose a hybrid spatial retrieval scheme. As shown in Figure 2 (Dense Retrieval Module), this method combines sparse spatial relevance scores from the database with dense semantic similarity scores from text embeddings. This enables the system to rank retrieved spatial objects based on their spatial relevance to the input question, as detailed in Section 4.3. Third, given both spatial and semantic constraints, we formulate a multi-objective optimization problem to balance these factors. The system computes the Pareto front of candidate answers, and the LLM dynamically trades off among these solutions to generate an optimal response. This step is covered in Section 4.4.
@@ 4.2 Sparse Spatial Retrieval
The answer to a spatial reasoning question must meet specific spatial constraints. The spatial candidate set $C_{s}(q)$ consists of all possible answers $y$ that satisfy a set of spatial constraints $\mathcal{C}_{s}(q)$. Formally, we define:
@@
Given the diversity and potentially complex nature of these constraints, Large Language Models (LLMs) often struggle to directly construct a complete and executable spatial query from user input. To bridge this gap, we structure the spatial query incrementally, allowing the LLM to systematically populate the required components.
Our approach follows three key steps: 1) Geometry Recognition: Identify and extract the reference spatial objects $G_{r}$ and candidate target spatial objects $G_{t}$ from the user’s input and extract their spatial footprints – geometries. 2) Query Function Selection: Determine the appropriate spatial function $\mathcal{F}_{s}$ based on the intended spatial relationship (e.g., containment, proximity). 3) Parameter Estimation: Assign numerical constraints $\epsilon$ to ensure precise spatial filtering (e.g., buffer radius).
@@ 4.2.1 Geometry Recognition
  • Point: $\mathcal{G}_{\text{point}}=\{g\mid g\in\mathbb{R}^{2},\dim(g)=0\}$ This category includes single points and multipoints, representing locations with negligible area. Examples include stop signs, address points, and a user’s current location. In spatial databases, these entities are typically represented as the ’Point’ geometry type.
@@
The complexity of a spatial query depends on the types of spatial footprints of objects involved. For simpler queries, such as ”finding the nearest bus stop from a given location”, only point geometries are required, and the spatial candidate set is
@@
where $g_{\text{point}}\subseteq\mathcal{G}_{\text{point}}$ represents a point object (e.g., given location), $\epsilon$ is the distance threshold. For more complex queries, such as ”I will walk from home to the university campus along 7th Street and Jones Street; please recommend a café where I can buy breakfast on my walk.”, multiple geometry types must be considered, and the spatial candidate set is
@@ 4.2.2 Query Function Recognition and Parameter Estimation
4.3 Hybrid Spatial Objects Ranking
The spatial relevance score $f_{s}$ consists of two components: a score derived from sparse spatial retrieval from the spatial database and a score from spatial dense retrieval based on text similarity between the question and the spatial descriptions of candidate objects. Formally, we define:
$f_{s}=\lambda_{s}f_{s}^{\text{sparse}}+\lambda_{d}f_{s}^{\text{dense}},$ (8)
where $\lambda_{s}$ and $\lambda_{d}$ are weighting coefficients controlling the contribution of each score.

### v4

# Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Geospatial Reasoning Questions

@@ Abstract
Answering real-world geospatial questions—such as finding restaurants along a travel route or amenities near a landmark—requires reasoning over both geographic relationships and semantic user intent. However, Existing large language models (LLMs) lack spatial computing capabilities and access to up-to-date, ubiquitous real-world geospatial data, while traditional geospatial systems fall short in interpreting natural language. To bridge this gap, we introduce Spatial-RAG, a Retrieval-Augmented Generation (RAG) framework designed for geospatial question answering. Spatial-RAG integrates structured spatial databases with LLMs via a hybrid spatial retriever that combines sparse spatial filtering and dense semantic matching. It formulates the answering process as a multi-objective optimization over spatial and semantic relevance, identifying Pareto-optimal candidates and dynamically selecting the best response based on user intent. Experiments across multiple tourism and map-based QA datasets show that Spatial-RAG significantly improves accuracy, precision, and ranking performance over strong baselines.
@@ 1 Introduction
Figure 1: A real-world spatial reasoning question with nearby spatial objects. Areas that satisfy the spatial constraint are highlighted in purple.
Spatial reasoning questions are those that require spatial computing to resolve relationships between objects, positions, or movements in space. Extensive research has been conducted on abstract spatial reasoning tasks such as mental rotation, block manipulation, and robot navigation, which rely on simplified, small-scale, and often purely geometric representations, typically addressed using techniques from computer vision and robotics. In contrast, geospatial reasoning involves interpreting large-scale, real-world geographic data where spatial information is deeply entangled with rich semantics chen2014parameterized ; mai2021geographic ; kefalidis2024question . For example, urban routing decisions depend not only on road geometries but also on attributes such as traffic regulations, land use, and temporal constraints. Travel plan recommendations should not only consider minimizing the travel distance, but also maximize the quality of the attractions according to their descriptions and reviews xietravelplanner .
Geospatial reasoning has a longstanding role in AI research, yet classical methods—such as spatial databases and GIS query systems—lack the ability to effectively interpret users’ natural language questions mai2021geographic . On the other hand, large language models (LLMs) exhibit strong linguistic competence but struggle with spatial computing and geospatial grounding mai2024opportunities . Recent efforts to bridge this gap have focused on prompt engineering manvigeollm ; gurnee2024language , but these approaches heavily rely on LLMs’ internal knowledge, which remains limited in generalization and spatial reasoning capabilities, significantly suffering from geographic bias faisal2023geographic ; manvi2024large ; wu2024torchspatial , and being susceptible to obsolescence as knowledge evolves. Some work has explored fine-tuning LLMs on spatial tasks ji2023evaluating ; manvigeollm ; zhang2024bb , but the resulting models are often tailored to narrow applications, constrained datasets, or specific geographic domains. Therefore, there remains a critical need for a general-purpose geospatial reasoning framework that synergizes semantic understanding and spatial computation while ensuring access to real-world, vast, fast-changing, and complex geospatial data.
@@
To fill this gap, this paper aims to augment LLMs with capabilities of spatial reasoning and accessibility to real-world geospatial data. For example, as illustrated in Figure 1, answering the question requires LLMs to elicit and formulate the user’s textual request into the problem of “finding points near the polyline" and solve it based on a geospatial map (database) with semantic information (e.g., customer reviews and location profiles). Then, it also requires inferring user intent to select the spatially and semantically preferred candidates. Thus, the system must seamlessly integrate structured spatial retrieval with unstructured text-based reasoning, ensuring both spatial accuracy and contextual understanding. Specifically, we extend Retrieval-Augmented Generation (RAG) into geospatial information retrieval and reasoning, bridging the gap between structured spatial databases and unstructured textual reasoning. RAG has demonstrated its effectiveness in knowledge-intensive tasks, such as question answering (QA) siriwardhana2023improving , by retrieving domain-specific documents to enhance LLM responses. However, existing RAG systems primarily focus on retrieving and generating textual content and lack the spatial intelligence required for spatial reasoning tasks, especially tasks that involve understanding and computing complex spatial relationships among geometries, including points, polylines, and polygons.
@@
In this paper, we introduce Spatial Retrieval-Augmented Generation (Spatial-RAG), a new framework that unifies text-guided spatial retrieval with spatially aware text generation under multi-objective optimization scenario. Specifically, to identify spatially relevant candidate answers, we propose a novel spatial hybrid retrieval module synergizing spatial sparse and dense retrievers. To rank the candidates and generate the final answers, we propose to fuel the generator with retrieved results on the Pareto frontier based on a spatial and semantic joint ranking strategy. Our contributions are summarized as follows:
@@
  • A generic spatial RAG framework: We introduce spatial-RAG, the first framework that extends RAG to geospatial question answering, to tackle a broad spectrum of spatial reasoning tasks, such as geographic recommendation, spatially constrained search, and contextual route planning. Our approach seamlessly integrates spatial databases, LLMs, and retrieval-based augmentation, enabling effective handling of complex spatial reasoning questions directly within the familiar operational paradigm of LLMs.
@@
  • Sparse-dense spatial hybrid retriever: We propose a hybrid retrieval mechanism that combines spatial sparse retrieval (e.g., SQL-based structured queries) with spatial dense retrieval (e.g., LLM-powered semantic matching). This dual approach ensures that retrieved results align both spatially and semantically with the user’s query, synergizing spatial computing and geographical text understanding.
@@
  • Multi-objective guided spatial and semantic text generator: To handle both spatial constraints and semantic intents in the spatial question-answering task, we introduce a multi-objective optimization framework that dynamically balances trade-offs between spatial and semantic relevancy to the user’s query. This ensures that the generated responses are both geospatially accurate and linguistically coherent.
@@
  • Real-world evaluation: We evaluated our method on multiple real-world datasets consisting of user-generated QA pairs about various spatial entities. The experiments demonstrate the model’s ability to handle spatial reasoning questions grounded in real-world scenarios.
@@
Through these innovations, Spatial-RAG significantly enhances the spatial reasoning capabilities of LLMs, bridging the gap between structured spatial databases and natural language QA.
@@ 2.1 Retrieval augmented generation
Retrieval-Augmented Generation (RAG) is a hybrid approach that integrates retrieval systems and generative models to enhance factual accuracy and contextual relevance in natural language generation fan2024survey . Unlike conventional language models that rely solely on parametric memory, RAG dynamically retrieves relevant external knowledge before generating a response. In RAG lewis2020retrieval , a retrieval module fetches relevant passages from a large-scale knowledge corpus (e.g., Wikipedia), which are then fused with the question context to generate a more informed response. This technique has proven particularly effective in open-domain question answering (QA), fact verification, and context-aware text generation. RAG systems have expanded beyond text and document retrieval to incorporate a wide variety of data types he2024g — tables, graphs, charts, and diagrams. While RAG has been widely explored, its application in spatial reasoning question answering remains an unexplored research area. Existing studies have primarily focused on knowledge-grounded dialogues yu2024llms but often struggle with integrating spatial computation into the question-answering process effectively.
2.2 Geospatial question and answering
@@
Spatial questions in domain-specific applications can generally be categorized into two distinct types: 1) Textual knowledge-based spatial questions These are spatial questions that can be answered by traditional QA methods without the need for spatial computation and reasoning lietard2021language . For example, the question "What is the population of Los Angeles city?" falls under this category. Despite their spatial context, these questions are essentially text-based and, hence, can be effectively addressed using traditional Retrieval-Augmented Generation (RAG) methods christmann2024rag . 2) Spatial reasoning questions This category encapsulates spatial questions that demand a model’s capability to comprehend and reason with spatial data and spatial relationships. A common example is a model being presented with textual information describing the spatial relationships among multiple objects li2024advancing . An example question could be, "What is the position of object A relative to object B?", where objects A and B are locations or entities specified on the map. Resolving such queries requires a profound understanding of spatial concepts and robust reasoning skills, which largely depend on the model’s training to handle spatial data. Several studies (mai2024opportunities, ; roberts2023gpt4geo, ) have investigated the capacity of LLMs to understand spatial concepts, yet these models often struggle with accurate reasoning even after fine-tuning. Other research li2023geolm has attempted to enhance this ability by converting geolocation coordinates into addresses to enrich the semantic context. However, these improvements tend to be marginal and are mostly limited to straightforward reasoning tasks like describing positions. Moreover, many existing methods rely on predefined sets of actions tailored to specific tasks.
@@ 3 Problem formulation
In this study, our primary focus is Geospatial Reasoning Questions. We formulate the problem as follows: Given a query $q$, the system aims to generate an answer $y^{*}$, which maximizes the joint spatial and semantic scores while satisfying the spatial constraints in the query $q$. For example, in Figure 1, the desired answer is a restaurant that satisfies a spatial constraint—it must be within walking distance of the route. Additionally, it should ideally be located along the street (spatial score) and preferably offer waffle fries (semantic score). This problem can be formulated as the following multi-objective optimization problem:
@@
s.t. $y\in C_{s}(q),\quad y\in C_{k}(q),\;\lambda_{s}\geq 0,\lambda_{k}\geq 0,$ $\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1,$
@@
Note that this problem can have multiple valid solutions depending on the trade-off parameters $\lambda_{s}$ and $\lambda_{k}$, forming a Pareto frontier. Existing approaches are unable to solve this problem effectively, as it demands a synergistic capability in both geospatial and semantic reasoning. Specifically: 1) it requires accurately determining whether a candidate $y$ satisfies the spatial constraints expressed in the query $q$. 2) it involves ranking candidates based on both spatial and semantic relevance, which are interrelated yet provide complementary signals; and 3) it must ensure that no high-quality answers are overlooked under different trade-offs between spatial and semantic aspects. Current methods, which typically excel in either geospatial reasoning (e.g., spatial databases) or semantic reasoning (e.g., large language models and their variants), cannot address all of these requirements without substantial effort to integrate both capabilities seamlessly.
4 Spatial-RAG for geospatial reasoning questions
@@ 4.1 Overview
Our Spatial-RAG is illustrated in Figure 2, which consists of three key stages: First, to construct the spatial candidate set $C_{s}$, the system must precisely define spatial constraints and then retrieve spatial objects that satisfy them. As depicted in Figure 2 (a) Sparse spatial retrieval, we achieve this by parsing the input natural language questions into a spatial SQL query, which will be executed on the spatial database to efficiently retrieve relevant spatial objects from the database. This process is detailed in Section 4.2. Second, to effectively compute spatial relevance $f_{s}(q,y)$ while integrating textual information, we propose a hybrid spatial retrieval scheme. This method combines sparse spatial relevance scores from the database with dense semantic similarity scores from text embeddings. This enables the system to rank retrieved spatial objects based on their spatial relevance to the input question, as detailed in Section 4.3. Third, given both spatial and semantic scores, we formulate a multi-objective optimization problem to balance these factors. The system computes the Pareto frontier of candidate answers, and the LLM dynamically trades off among these solutions to generate an optimal response. This step is covered in Section 4.4.
@@ 4.2 Sparse spatial retrieval
The answer to a spatial reasoning question must meet specific spatial constraints. The spatial candidate set $C_{s}(q)$ consists of all possible answers $y$ that satisfy a set of spatial constraints $\mathcal{C}_{s}(q)$:
@@
Given the diversity and complex nature of these constraints, LLMs often struggle to directly construct a complete and executable spatial query from user input. To bridge this gap, we parse the spatial query incrementally, allowing LLMs to systematically populate the required components. Our approach follows three key steps: 1) Geometry recognition: Identify and extract the reference spatial objects $G_{r}$ and candidate target spatial objects $G_{t}$ from the user’s input and extract their spatial footprints – geometries. 2) Query function selection: Determine the appropriate spatial function $\mathcal{F}_{s}$ based on the intended spatial relationship (e.g., containment, proximity). 3) Parameter estimation: Assign numerical constraints $\epsilon$ to ensure precise spatial filtering (e.g., buffer radius).
@@ 4.2.1 Geometry recognition
  • Point: $\mathcal{G}_{\text{point}}=\{g\mid g\in\mathbb{R}^{2},\dim(g)=0\}$ This category includes single points and multipoints, representing locations with negligible area. Examples include stop signs, points of interest, and a user’s current location. In spatial databases, these entities are typically represented as the ’Point’ geometry type.
@@
The complexity of a spatial query depends on the types of spatial footprints of objects involved. For simpler queries, such as "finding the nearest bus stop from a given location", only point geometries are required, and the spatial candidate set is
@@
where $g_{\text{point}}\subseteq\mathcal{G}_{\text{point}}$ represents a point object (e.g., given location), $\epsilon$ is the distance threshold. For more complex queries, such as "I will walk from home to the university campus along 7th Street and Jones Street; please recommend a café where I can buy breakfast on my walk.", multiple geometry types must be considered, and the spatial candidate set is
@@ 4.2.2 Query function recognition and parameter estimation
4.3 Hybrid spatial relevance scoring
The spatial relevance score $f_{s}$ consists of two components: a score derived from sparse spatial retrieval from the spatial database and a score from dense spatial retrieval based on text similarity between the question and the spatial descriptions of candidate objects. Formally, we define:
$f_{s}=\lambda_{p}f_{s}^{\text{sparse}}+\lambda_{d}f_{s}^{\text{dense}},$ (8)
where $\lambda_{p}$ and $\lambda_{d}$ are weighting coefficients controlling the contribution of each score.